A coin is biased such that it is 1:2 for heads. If the coin is tossed 6 times, find the probabilities of obtaining various number of heads. 


1
Expert's answer
2022-03-08T09:21:01-0500

The probability that we get head "p = \frac{2}{3}," that we get tail "q = 1-p=1-\frac{2}{3}=\frac{1}{3}."

Obviously the number of heads X we get after 6 tosses may be any value of 0, 1, 2, 3, 4, 5, 6.

We have a Bernoulli trial - exactly two possible outcomes, "success" (we get head) and "failure" (tail) and the probability of success is the same every time the experiment is conducted (the coin is tossed).

The pobability of each result

"P(X=k)=\begin{pmatrix}
n \\
 k
\end{pmatrix}\cdot p^k \cdot q^{n-k}=\begin{pmatrix}
6 \\
 k
\end{pmatrix}\cdot \begin{pmatrix}
 \cfrac{2}{3}
\end{pmatrix}^k
\cdot \begin{pmatrix}
 \cfrac{1}{3}
\end{pmatrix}^{n-k}="

"=\cfrac{n!}{k!\cdot{(n-k)!}} \cdot \begin{pmatrix}
 \cfrac{2}{3}
\end{pmatrix}^k
\cdot \begin{pmatrix}
 \cfrac{1}{3}
\end{pmatrix}^{n-k};"


"P(X=0)=\cfrac{6!}{0!\cdot{6!}} \cdot \begin{pmatrix}
 \cfrac{2}{3}
\end{pmatrix}^0
\cdot \begin{pmatrix}
 \cfrac{1}{3}
\end{pmatrix}^{6} =\cfrac{1}{3^6}=\cfrac{1}{729};"


"P(X=1)=\cfrac{6!}{1!\cdot{5!}} \cdot \begin{pmatrix}
 \cfrac{2}{3}
\end{pmatrix}^1
\cdot \begin{pmatrix}
 \cfrac{1}{3}
\end{pmatrix}^{5} =\cfrac{6\cdot2}{3^6}=\cfrac{12}{729};"


"P(X=2)=\cfrac{6!}{2!\cdot{4!}} \cdot \begin{pmatrix}
 \cfrac{2}{3}
\end{pmatrix}^2
\cdot \begin{pmatrix}
 \cfrac{1}{3}
\end{pmatrix}^{4} =\cfrac{5\cdot6\cdot2^2}{2\cdot3^6}=\cfrac{60}{729};"


"P(X=3)=\cfrac{6!}{3!\cdot{3!}} \cdot \begin{pmatrix}
 \cfrac{2}{3}
\end{pmatrix}^3
\cdot \begin{pmatrix}
 \cfrac{1}{3}
\end{pmatrix}^{3} =\cfrac{4\cdot5\cdot6\cdot2^3}{2\cdot3\cdot3^6}=\cfrac{160}{729};"


"P(X=4)=\cfrac{6!}{4!\cdot{2!}} \cdot \begin{pmatrix}
 \cfrac{2}{3}
\end{pmatrix}^4
\cdot \begin{pmatrix}
 \cfrac{1}{3}
\end{pmatrix}^{2} =\cfrac{5\cdot6\cdot2^4}{2\cdot3^6}=\cfrac{240}{729};"


"P(X=5)=\cfrac{6!}{5!\cdot{1!}} \cdot \begin{pmatrix}
 \cfrac{2}{3}
\end{pmatrix}^5
\cdot \begin{pmatrix}
 \cfrac{1}{3}
\end{pmatrix}^{1} =\cfrac{6\cdot2^5}{3^6}=\cfrac{192}{729};"


"P(X=6)=\cfrac{6!}{6!\cdot{0!}} \cdot \begin{pmatrix}
 \cfrac{2}{3}
\end{pmatrix}^6
\cdot \begin{pmatrix}
 \cfrac{1}{3}
\end{pmatrix}^{0} =\cfrac{2^6}{3^6}=\cfrac{64}{729}."